# SLIM: Sim-to-Real Legged Instructive Manipulation via Long-Horizon Visuomotor Learning

### v1

@@ D-A Control Latency Management
    Asynchronous low-high inference: most of the model latency comes from the high-level inference, which takes around 20 ms, but can sometimes peak to 40 ms. This could delay the low-level policy (50HZ) considerably and the robot to falter, if the two inferences are performed in sequence. Alternatively, in the asynchronous mode, the low-level policy will reuse the previous locomotion command if the new one has not arrived yet, and thus still maintain its 50HZ frequency. This leads to a more stable gait of the quadruped.
@@
    Sensor delay randomizations: motor sensor reading delay is sampled from $(3,13)$ms per episode with $\pm$2ms variation uniformly per control step. Camera observation delay is sampled from $(30,50)$ms uniformly per step with observations taken at 60HZ. Low-level control delay is sampled from $(5,7)$ms uniformly per step. High-level control delay should have been sampled from $(20,40)$ms, but was accidentally set to none during implementation. We did not repeat the full evaluation after discovering this bug, due to limited time. We did test introducing an additional 40ms control delay during physical deployment to the existing high-level policies. It did not impact performance in any significant way. We do plan to train models with the correct delay randomization in the future.
@@
D-B Safety Measures for the Physical Robot
@@
Appendix E Real-World Experiment Details
E-A Containers and Graspable Objects
@@
E-B Real-world Evaluation Protocols
@@
TABLE V: Scene Setups for Real-World Quantitative Evaluations. Legends: $\uparrow$ denotes the position robot with its orientation aligned with the direction of $\uparrow$. Uppercase letters denote baskets and lower case letters denote the objects to be grasped. The color of an object is denoted both by the background color and the character (e.g. g denotes the green cube and R denotes the red basket.)
@@
    To ensure the repeatability of the scene setup, we use the Standard spatial layout as shown in Figure 8, as it is easy to ensure accuracy of the scene setup in terms of spatial positions. In practice, we mark the positions for the robot at the center of the square. Similarly, we mark the 4 corner positions on the lobby floor for placing the objects and baskets. The size of the square is about $3\times 3$ meters.
@@
    We follow a scene initialization procedure as detailed in Table V across the 20 episodes to ensure good coverage of the relative positioning of the objects, the target object and basket in the language instruction, and the colors of the objects.

### v2

@@ E-A Control Latency Management
    Asynchronous low-high inference: most of the model latency comes from the high-level inference, which takes around 20 ms, but can sometimes peak to 40 ms. This could delay the low-level policy (50HZ) considerably and make the robot falter, if the two levels are executed in sequence. Alternatively, in the asynchronous mode, the low-level policy will reuse the previous locomotion command if the new one has not arrived yet, and thus still maintain its 50HZ frequency. This leads to a more stable gait of the quadruped.
@@
    Sensor delay randomizations: motor sensor reading delay is sampled from $(3,13)$ms per episode with $\pm$2ms variation uniformly per control step. Camera observation delay is sampled from $(30,50)$ms uniformly per step with observations taken at 60HZ. Low-level control delay is sampled from $(5,7)$ms uniformly per step. About the high-level inference delay, we evaluated the policies trained without high-level inference delay and compared the performances with and without test time high-level inference delays. We did not find clear performance differences in terms of the success rate. Therefore, we did not use high-level control delay in policy training in this work. Meanwhile, it is an option that is provided in our implementation and can be turned on whenever necessary.
@@
E-B Safety Measures for the Physical Robot
@@
Appendix F Real-World Experiment Details
F-A Containers and Graspable Objects
@@
F-B Real-world Evaluation Protocols
@@
TABLE VI: Scene Setups for Real-World Quantitative Evaluations. Legends: $\uparrow$ denotes the position robot with its orientation aligned with the direction of $\uparrow$. Uppercase letters denote baskets and lower case letters denote the objects to be grasped. The color of an object is denoted both by the background color and the character (e.g. g denotes the green cube and R denotes the red basket.)
@@
    To ensure the repeatability of the scene setup, we use the Standard spatial layout as shown in Figure 8, as it is easy to ensure accuracy of the scene setup in terms of spatial positions. In practice, we mark the positions for the robot at the center of the square. Similarly, we mark the 4 corner positions on the lobby floor for placing the objects and baskets. The size of the square is about $2\times 2$ meters.
@@
    We follow a scene initialization procedure as detailed in Table VI across the 20 episodes to ensure good coverage of the relative positioning of the objects, the target object and basket in the language instruction, and the colors of the objects.
F-C Visualization Predicted Segmentation and Depth Maps
[…]
Figure A.17: Maps prediction results for some key frames of a deployment session. Time flows from top to bottom, and each row corresponds to one time step. From left to right, the columns represent the RGB frame, the predicted depth map, the predicted segmentation mask, and our added note for better interpretation of different stages.
In Figure A.17, we show the robot’s prediction results for depth and object segmentation, during an arbitrary outdoor deployment session. We emphasize again that the robot was deployed in this zero-shot scenario never modeled by our simulator, and no model finetuning was performed. Despite this, we can see that the prediction results are high-quality, given only a very short history ($0.5$ seconds) of a single wrist-mounted RGB camera and the robot proprioceptive state. While searching for the target cube, the perception model had to locate it with just a handful of pixels (first row). Right after the gripper holding and lifting the cube above the ground, the model correctly predicted that the target had switched from the cube to a basket (cube’s green color turned red; third row). Moreover, while searching for the target basket in the outdoor environment, the cluttered background did not confuse the depth and segmentation prediction (fourth row). The perception model was also able to correctly predict the segmentation mask for the robot itself (gripper fingers and quadruped in orange; fifth row). Overall, we find that in practice, this high-quality visual information bottleneck not only greatly improves the generalization ability of our RL policy, but also facilitates our debugging of the entire robotic system as it separates the errors of the perception model from those of the policy.